# Kishu: Time-Traveling for Computational Notebooks

### v1

@@ 7.1. Experiment Setup
We select a total of 8 data science notebooks from Kaggle Grandmaster-level users or tutorials of established tools (e.g., Ray). LABEL:tbl:workload reports our selected notebooks’ dataset sizes and runtimes.
Additionally, we select 146 data science library classes commonly used in data science notebooks, on which we evaluate the robustness of Kishu’s modification detection and time-traveling correctness. LABEL:tbl:library reports a classification of the libraries.
@@ Methods
  • ElasticNotebook (Li et al., [n.d.]): An application-level notebook migration tool that balances data serialization and cell recomputation to achieve optimized session replication times.
@@
We additionally compare the overhead of Kishu’s update detection mechanism with the following methods:
@@
  • Kishu (Check all): Always perform update detection for all Co-variables in the session state after each cell execution, regardless of whether they were accessed by the user in the previous cell.
We consider these methods regarding session state delta tracking time (§ 7.6) to gauge the effectiveness of our overhead reduction via identifying candidates for updated Co-variables.
Environment
Experiments are performed on an Ubuntu server with 2 AMD EPYC 7552 48-Core Processors and 1TB RAM. The checkpoints for all methods are written to a mounted NFS, with disk read and write speeds of 519.8 MB/s and 358.9 MB/s, respectively.
Time measurement
We measure (1) (incremental) checkpoint time as time spent checkpointing (including both tracking and data writing) after each cell execution, (2) checkout time as time to restore the state from checkpoint files, and (3) tracking overhead as time spent after each cell execution tracking Co-variable granularity updates. We clear the page cache between runs.
Reproducibility
Our implementation of Kishu, experiment notebooks, full list of data science library classes we test with, and scripts can be found in our Github repository. [^4] https://anonymous.4open.science/r/kishu-vldb-E724
@@ 7.2. Generalized and Robust Time Traveling
This section compares the robustness of Kishu’s time-traveling to existing methods. We attempt to checkpoint and checkout session states containing objects from the 146 data science library classes, and compare the number of classes each method fails to checkout. We report the results in LABEL:fig:experiment_robust. Kishu completes checkpoint and checkout for all 146 libraries, notably handling 6 classes involving multiprocessing and/or off-CPU data and 7 unserializable classes that CRIU and DumpSession fail to checkpoint and/or checkout, respectively. This is because (1) unlike CRIU, Kishu relies on object reductions to store Co-variables, hence it can store distributed or off-CPU data (e.g., Ray’s dataset(Team, 2024b) or on-GPU tensors(TensorFlow, 2024b; PyTorch, 2024)) and (2) unlike DumpSession, Kishu utilizes fallback recomputation, allowing it to restore Co-variables containing (1) unserializable objects (e.g., pl.LazyFrame (Polars, 2024b)) or (2) serializable objects that can’t deserialize (e.g., bokeh.figure (Bokeh, 2024)). We present a summary of these noteworthy object classes in LABEL:tbl:experiment_robust_example_table.
@@ 7.2.1. Accurate Delta Detection
This section verifies accuracy of Kishu’s delta detection. For each class, we test whether two VarGraphs generated for a class object differ before and after (1) updating a class attribute (e.g., model.key = ’A’) or (2) updating nothing. We count the number of VarGraph differences for case (1) as successes and case (2) as false positives.
We report the results in LABEL:tbl:experiment_fast_capture_accuracy. Kishu’s delta detection accurately captures object updates in 120 classes. While Kishu reports false positives in 14 classes, (e.g., due to dynamically generated reachable objects), these false positives only affect Kishu’s efficiency (i.e., during checkpointing/checkout); however, Kishu maintains accuracy by considering these objects to be updated on access. We also find that 12 classes contain silent pickling errors (§ 6.2); nevertheless, Kishu reports these objects to be updated on access similar to false positives, and users may force their (fallback) recomputation if needed (§ 6.2). Notably, Kishu does not have false negatives: Kishu will always report if an object is changed.
@@ 7.3. Small Incremental Checkpoint Sizes
We report the results in LABEL:fig:experiment_cheap_checkpoint. Kishu’s cumulative checkpoint size is consistently the smallest on all notebooks and is up to 4.55$\times$ smaller than the next best alternative (HW-LM). ElasticNotebook, while the next best method on 6/8 notebooks and similarly has fault-tolerant mechanisms to successfully checkpoint all 8 notebooks, can fall short in checkpointing time (§ 4.3). CRIU-Incremental, despite also incrementally checkpointing, is not the next best method on any notebook, losing to ElasticNotebook and DumpSession on 6 and failing to checkpoint on 2 as it (1) incrementally checkpoints at the coarser memory page level (§ 2.3), and (2) does not handle off-CPU data and multiprocessing (§ 7.2). DumpSession fails on the Qiskit notebook due to its inability to handle unserializable data, and checkpointing with CRIU incurs prohibitive storage costs (94GB on TPS) as it non-incrementally checkpoints at the OS-level.
@@ 7.4. Low Incremental Checkpoint Time
We report the results in LABEL:fig:experiment_fast_checkpoint. Kishu’s cumulative checkpointing time is the lowest on 5/8 notebooks, being only up to 15.5% of notebook runtime (HW-LM) and up to 5.12$\times$ faster (HW-LM) than the next best alternative on these notebooks. While CRIU-Incremental checkpoints faster compared to Kishu on 3/8 notebooks owing to writing memory pages being faster than serialization for unit amount of data, the difference is negligible (up to 3.03$\times$, StoreSales) compared to the reliability issues (§ 7.2), space inefficiency (§ 7.3), and it being consistently slower than Kishu by more than an order of magnitude for checkout (§ 7.5). Compared to ElasticNotebook, Kishu achieves fast checkpointing by being EAFP-based (Goodger, 2014): if it fails to store a Co-variable, it will simply recompute it upon checkout via fallback recomputation. This allows it to skip the per-execution profiling steps (i.e., for data sizes and serializability) present in ElasticNotebook’s LBYL-based (Goodger, 2014) optimization (for which data to store/recompute), which can cause checkpoint times slower than DumpSession on 2/8 notebooks.
@@ 7.5.1. Fast Execution Undo
For each notebook, we measure the time it takes to undo cells containing dataframe operations (notebook-D) and plot modifications (notebook-P).
We report the results in LABEL:fig:experiment_checkout_undo. Kishu achieves sub-second cell execution rollbacks on all test cases, and is up to 8.18$\times$ faster than the next best alternative (StoreSales-D). While CRIU-Incremental achieves checkpoint times comparable with Kishu, it is up to 36$\times$ slower for checking out (StoreSales-D) and the slowest method for undos on 5/6 notebooks, due to it needing to piece together the memory snapshot of the notebook process to restore from multiple (incremental) checkpoint files. Compared to CRIU, DumpSession, and ElasticNotebook, Kishu is the only method that consistently performs sub-second checkouts, showcasing the importance of performing incremental checkout (i.e., vs. non-incremental) by identifying state differences. For example, in the Sklearn notebook, the test case (cell 28) drops a column in an auxiliary dataframe that is 1.4MB in size (compared to the 133MB main dataframe). Kishu identifies that it only needs to load the auxiliary dataframe from before the cell execution and undoes the operation in 0.4 seconds; however, other methods all require the entire session state to be overwritten with a complete load of checkpoint data, taking an upwards of 6 seconds to do so (and in CRIU and CRIU-Incremental’s case, also killing and restarting the current notebook process).
@@ 7.5.2. Fast Path Exploration
We report the results in LABEL:fig:experiment_checkout_branch. Similar to § 7.5.1, Kishu performs sub-second branch switching on 4/6 notebooks by updating (only) the models and plots differing between the branches (i.e., not the input dataframes) and does so up to 9.02$\times$ faster than the next best alternative (TorchGPU). In the StoreSales notebook, even when there is considerable divergence between the branches (i.e., new auxiliary dataframes are created along with ML models and plots), Kishu still performs branch switching at a relatively fast 1.73 seconds, which is 4.19$\times$ faster than the next method (DumpSession).
@@ 7.6. Fast Delta Detection
This section investigates overhead of Kishu’s Co-variable granularity state tracking. We compare the cumulative time Kishu spends to track per-execution state delta with Kishu (Check all) and IPyFlow.
We report the results in LABEL:fig:experiment_fast_capture. Kishu is consistently the fastest method for detecting the state delta and is up to 11.42$\times$ faster than the next best method out of IPyFlow and Kishu (Check all). Notably, the detection overhead is a maximum of 3.0% of the notebook runtime (Sklearn), which is (only) an average of 0.23 seconds per cell. Kishu’s fast delta detection can be attributed to it operating at a coarser-grained granularity uniquely suitable for Kishu’s use case compared to IPyFlow’s symbol-level detection (§ 2.3), which can incur up to 2.30$\times$ overhead in terms of notebook runtime (HW-LM). Finally, as shown by the comparison between Kishu and Kishu (Check all), Kishu’s ability to limit the number of objects it has to check for changes after each cell execution via identifying candidate Co-variables for updates is necessary: it reduces the overhead from 136s to 10s on the Sklearn notebook (University, 2021), which we study in § 7.6.1.
7.6.1. Usefulness of Candidates
This section verifies usefulness of Kishu’s identification of candidate Co-variable updates by studying the number of variables (i.e., names, e.g. x) and percentage of session state (w.r.t. memory size) that Kishu checks for updates after each cell execution on the Sklearn notebook.
We report the results in LABEL:fig:experiment_fast_capture_candidates: we observe that while the user continually defines new variables throughout the session, each cell execution is highly atomic and refers to only a few variables (0-8) at a time (LABEL:fig:experiment_fast_capture_candidates_vars). Kishu successfully captures the references, which it utilizes to then check only a small portion of objects ($\sim$15%, LABEL:fig:experiment_fast_capture_candidates_objs). This effectively bounds Kishu’s update detection overhead to be largely independent of the memory size of the session state.
@@ 8. Related Work
This section covers related work in Time-Traveling in Databases and conventional programs, other applicable methods for checkpointing/restoring notebook states, and notebook lineage tracing.
Time-Traveling Databases
Many existing DBMS support Point-in-time-Recovery (PITR) by using a combination of (incremental) checkpointing and logging (Oracle, 2024a; Group, 1996b; Soroush and Balazinska, 2013; Schule et al., 2019; Xu et al., 2017; Antonopoulos et al., 2019). Incremental checkpoints are commonly created at the row level where deltas would store differing records between versions of tables (Antonopoulos et al., 2019; Oracle, 2024a), while individual operations would be recorded in logs using techniques such as ARIES (Mohan et al., 1992; Mohan and Rothermel, 1988) or WBL (Arulraj et al., 2016). Then, to return to a previous state (e.g., in the case of a database failure), the DBMS will load the appropriate checkpoint and replay the logs up to the desired state (e.g., defined by state snapshots as in MVCC (Bernstein and Goodman, 1983)). More recently, PITR features have been added to general object stores such as SQLite (LITEREPLICA, 2024), MLFlow (Zaharia et al., 2018), and DeltaLake (Armbrust et al., 2020), which can hold arbitrary objects (i.e., blobs) in addition to conventional tables. PITR for DBMS or object stores is largely orthogonal and inapplicable to time-traveling in notebooks, as unlike (full) PITR after DBMS failure, notebook state restoration needs to be performed into an existing kernel at interactive speeds. Hence, time-traveling (i.e., restoring) via operation replay is less desirable for notebooks: while notebook cell executions are typically fewer in count (Google and X, 2022) compared to row transactions, each operation can take much longer (e.g., training an ML model vs. writing a row). Kishu addresses these differences in problem setting by incrementally checkpointing the state at every step to ensure interactive state restoration times while utilizing novel delta detection techniques to maintain low checkpoint storage overhead, only utilizing (logged) cell replay for fallback recomputations to restore missing or problematic data.

### v2

@@ 7.1. Experiment Setup
We select 8 data science notebooks from Kaggle Grandmaster-level users or Github-hosted tool tutorials (e.g., Ray) (LABEL:tbl:workload), each featuring a popular data science library, which we categorize as in-progress (3/8) or final (5/8), with the former containing out-of-order cell executions and the latter lacking them. Notably, the final notebooks are also cleaned and contain memory/runtime optimizations. We empirically verify all notebooks follow traits discussed in § 2.2, and provide more details on our categorization (and its implications) in the appendix (§ 10).
We also select 146 data science library classes (summarized in LABEL:tbl:library) commonly used in data science notebooks, on which we evaluate Kishu’s correctness and robustness.
@@ Methods
  • ElasticNotebook (Li et al., 2023a): An application-level notebook migration tool that balances data serialization and cell recomputation to achieve optimized session replication times.
  • Kishu+Det-replay: A checkpoint-optimized version of Kishu based on operation replay (Mohan et al., 1992; Netezza, 2024) that uses manual annotation [^6] Automatically detecting non-/determinism in executions is out of scope for this work. to skip incremental checkpointing for deterministic cells. These deterministic cells are replayed as necessary on checkout.
@@
We additionally compare the overhead of Kishu’s update detection mechanism with these tracking methods (§ 7.6):
@@
  • AblatedKishu (Check all): Always perform update detection for all Co-variables in the session state after each cell execution, regardless of whether they were accessed in the previous cell.
Methodology and Measurement
We run notebook cells sequentially and checkpoint after each cell execution. We checkout into the same state for Kishu and Kishu+Det-replay and into a fresh kernel/process for other methods. We measure the (1) checkpoint time (including both tracking and data writing) after each cell execution, (2) checkout time to restore the state from checkpoint files, and (3) tracking overhead spent by Kishu after each cell execution to track updates. We clear the page cache between runs.
Environment and Reproducibility
Experiments are performed on an Ubuntu server with 2 AMD EPYC 7552 48-Core Processors and 1TB RAM. All checkpoints are written to a mounted NFS, with disk read and write speeds of 519.8 MB/s and 358.9 MB/s, respectively. Our Github repository [^7] https://github.com/illinoisdata/kishu-vldb contains our implementation of Kishu, experiment notebooks/library classes, and scripts.
@@ 7.2. Generalized and Robust Time Traveling
This section compares the robustness of Kishu’s time-traveling to existing methods. We attempt to checkpoint and checkout session states containing objects from the 146 data science library classes and compare number of classes each method fails to checkout.
We report results in LABEL:fig:experiment_robust. Kishu completes time-traveling for all 146 libraries, handling 6 classes with multiprocessing and/or off-CPU data and 7 unserializable classes that CRIU and DumpSession fail on, respectively: unlike CRIU, Kishu utilizes reductions (§ 6.1) to store Co-variables, hence it can store distributed or off-CPU data (e.g., Ray’s dataset(Team, 2024b) or on-GPU tensors(TensorFlow, 2024b; PyTorch, 2024)) and unlike DumpSession, Kishu’s fallback recomputation allows it to restore Co-variables with (1) unserializable objects (e.g., pl.LazyFrame(Polars, 2024b)) or (2) serializable objects that can’t deserialize (e.g., bokeh.figure(Bokeh, 2024)). LABEL:tbl:experiment_robust_example_table summarizes these noteworthy classes.
@@ 7.2.1. Accurate Delta Detection
We verify Kishu’s delta detection accuracy by comparing two VarGraphs generated for each class object before and after (1) updating a class attribute (e.g., model.key = ’A’) or (2) updating nothing. We count the number of VarGraph differences for case (1) as successes and case (2) as false positives.
We report results in LABEL:tbl:experiment_fast_capture_accuracy. Kishu’s VarGraphs accurately captures object updates in 120 classes. While Kishu reports false positives in 14 classes, (e.g., due to dynamically generated reachable objects), they only affect Kishu’s efficiency (i.e., during time-traveling); however, Kishu maintains accuracy by considering these objects to be updated on access. We also find that 12 classes contain silent pickling errors (§ 6.2); nevertheless, Kishu reports these objects to be updated on access similar to false positives, and users may force their (fallback) recomputation if needed (§ 6.2). Notably, Kishu has no false negatives: Kishu will always report if an object is changed.
@@ 7.3. Small Incremental Checkpoint Sizes
We report results in LABEL:fig:experiment_cheap_checkpoint. Kishu’s cumulative checkpoint size is consistently the smallest (expectedly except for Kishu+Det-replay, explained shortly) and is up to 4.55$\times$ smaller than the next best alternative (HW-LM). ElasticNotebook, while the next best method on 6/8 notebooks and also has fault-tolerant mechanisms to checkpoint all 8 notebooks, can fall short in checkpointing time (§ 4.3). CRIU-Incremental, while also incrementally checkpointing, is not the next best method on any notebook, losing to ElasticNotebook and DumpSession on 6 and failing to checkpoint on 2 as it (1) incrementally checkpoints at the coarser memory page level (§ 2.4), and (2) does not handle off-CPU data and multiprocessing (§ 7.2). DumpSession fails on Qiskit as it cannot handle unserializable data, and CRIU incurs prohibitive storage costs (94GB, TPS) as it non-incrementally checkpoints at OS-level. While Kishu+Det-replay can save checkpoint storage cost of up to 3.95$\times$ vs. Kishu (StoreSales) by skipping checkpointing after deterministic cells, it (1) needs manual annotation and (2) can result in unacceptable checkout times (§ 7.5).
@@ 7.4. Low Incremental Checkpoint Time
We report results in LABEL:fig:experiment_fast_checkpoint. Kishu’s cumulative checkpointing time is the lowest (except for Kishu+Det-replay) on 5/8 notebooks, being only up to 15.5% of notebook runtime (HW-LM) and up to 5.12$\times$ faster (HW-LM) than the next best alternative on these notebooks. While CRIU-Incremental checkpoints faster than Kishu on 3/8 notebooks owing to memory dumping being faster than serialization for unit data, the difference is negligible (up to 3.03$\times$, StoreSales) compared to the reliability issues (§ 7.2), space inefficiency (§ 7.3), and slow checkout times (§ 7.5). Compared to ElasticNotebook, Kishu’s checkpointing is EAFP-based (Goodger, 2014): if it fails to store a Co-variable, it will simply recompute it upon checkout via fallback recomputation. This allows it to skip the profiling (i.e., for data sizes and serializability) required for ElasticNotebook’s optimization (for what to store/recompute), which causes checkpoint times slower than DumpSession on 2/8 notebooks.
@@ 7.5.1. Fast Execution Undo
For each notebook, we measure the time it takes to undo various dataframe and plot operation cells.
We report the results in LABEL:fig:experiment_checkout_undo. Kishu is the only method capable of incrementally checking out: it achieves sub-second cell execution rollbacks on all test cases, and is up to 8.18$\times$ faster than the next best alternative (StoreSales). While CRIU-Incremental achieves checkpoint times comparable with Kishu, it is up to 36$\times$ slower for checking out (StoreSales) and the slowest method for undos on 5/6 notebooks, due to it needing to piece together the memory snapshot of the notebook process to restore from multiple (incremental) checkpoint files. CRIU, DumpSession, and ElasticNotebook cannot incrementally checkout hence they cannot consistently perform sub-second undos. For example, the Sklearn notebook test case drops a column in an auxiliary dataframe that is 1.4MB in size (vs. the 133MB main dataframe). Kishu identifies that it only needs to load the auxiliary dataframe from before the cell execution and undoes the operation in 0.4 seconds; however, other non-Kishu+Det-replay methods all require the entire session state to be overwritten with a complete load of checkpoint data, taking an upwards of 6 seconds to do so (and in CRIU and CRIU-Incremental’s case, also killing and restarting the current notebook process).
@@ 7.5.2. Fast Path Exploration
We report the results in LABEL:fig:experiment_checkout_branch. Similar to § 7.5.1, Kishu performs sub-second branch switching on 4/6 notebooks by updating (only) models and plots differing between branches (i.e., not the input dataframes) and does so up to 4.18$\times$ faster than the next best alternative (StoreSales). While there is considerable divergence between branches in the StoreSales test case (i.e., new auxiliary dataframes are created along ML models and plots), Kishu still performs branch switching at a fast 1.73 seconds, which is 4.18$\times$ faster than the next method (DumpSession). While Kishu+Det-replay can potentially be faster than Kishu (TorchGPU) by replaying cells that allow it to bypass expensive data loading, it can also cause unacceptable checkout times (1050s, Cluster, from replaying an entire deterministic model fitting sequence); hence, cost-based optimization is required for Kishu+Det-replay to function, which we leave to future work.
@@ 7.6. Fast Delta Detection
This section investigates Kishu’s Co-variable granularity state tracking overhead by comparing the time taken by Kishu to track per-cell execution state delta with other tracking methods.
Cumulative Tracking Overhead (LABEL:tbl:experiment_fast_capture)
Kishu is consistently fastest at detecting state delta and is (1) up to 11.42$\times$ faster than the best out of IPyFlow and AblatedKishu (Check all) (HW-LM), and (2) only up to maximum of 2.03% of notebook runtime (Sklearn).
Per-Cell Tracking Overhead (LABEL:fig:experiment_fast_capture_percentage)
[…]
Notably, there exists further optimization opportunities for Kishu such as for (1) cells updating Co-variables with nested VarGraphs (e.g. list of strings text_neg, Sklearn cell 41) and (2) read-only printing cells (e.g., y_train[:10], HW-LM cell 67). Kishu incurs significant overhead on these cases (260$\times$ and 1.06$\times$ respectively; while they are of low absolute value due to the short cell runtimes (547ms/2ms and 2ms/2ms respectively), this indicates need for more efficient delta detection methods such as list hashing and rule-based detection, respectively (§ 6.2), which we leave for future exploration.
7.7. Workload Study
This section studies Kishu’s performance versus parameter sweeps on the degree of shared referencing between variables (i.e., Co-Variable size) (§ 7.7.1) and number of cell executions (§ 7.7.2).
7.7.1. Performance vs. Shared Referencing
We insert ten 64MB numpy arrays into a list (of state data in a Co-Variable). We evaluate Kishu’s checkpoint/checkout costs on cells modifying only one array in the list/Co-Variable. DumpSession’s and CRIU-Incremental’s performances are provided as comparison.
We report results in LABEL:fig:experiment_covariable. As Kishu detects updates and checkpoints at the Co-variable granularity, it checks for updates and then checkpoints all arrays in the list after each test cell. Kishu performs best when the list contains a low % of data in the state, as it can (1) limit the scope of update checking and (2) time-travel by saving/loading a small amount of state data. Kishu’s performance drops as the list Co-variable bundles the changed array with more unchanged data and is equivalent to DumpSession  [^9] Different serialization libraries cause different checkpoint/checkout time cost; there is also overhead with managing data with blob storage vs. writing directly to file (§ 6.1). when all 10 arrays (i.e., all data in the state) form one large Co-Variable: Kishu has to (1) check the whole state for updates and (2) save/update all data for checkpoint/checkout, while CRIU-Incremental can still checkpoint only the one changed array in the (640MB) Co-variable.
However, as our evaluation workloads suggest (LABEL:tbl:experiment_covariable_count), states typically consist of a large number of small Co-Variables (each containing 2.57% of state data on average, black vertical lines in LABEL:fig:experiment_covariable). For these typical cases, our Co-Variable-based approach significantly outperforms other baselines (§ 7.3, § 7.4).
7.7.2. Scalability to Long Notebook Sessions
We choose 2 Data Visualization notebooks (HW-LM and Qiskit, and randomly re-execute up to 1000 cells. [^10] This is the length of the longest observed notebook on Kaggle (Google and X, 2022), and $\sim$10 times the 97th percentile cell execution count of a notebook workload (Rule et al., 2018). We measure the (1) Checkpoint Graph size and (2) time to compute state difference for undoing 0-1000 cells (i.e., to a prior state) from the state after the 1000th cell execution.
We report results in LABEL:fig:experiment_scalability. The size of Kishu’s Checkpoint Graph scales linearly with the number of executed cells and is up to only 9MB at 1000 cells (LABEL:fig:experiment_scalability_a, HW-LM). Kishu’s time to compute state difference between current and target states scales linearly with the sum of cell count in the two states, up to 81ms for any checkout operation in a 1000 cell execution-long session. [^11] This is because Kishu traverses the Checkpoint Graph to find the lowest common ancestor state on checkout with an off-the-shelf algorithm (Wikipedia, 2024). These overheads are negligible on a more typical notebook—only 133KB and 5.8ms, respectively, on our longest experiment notebook (Qiskit, 85 cells).
@@ 8. Related Work
This section covers related work in PITR/checkpointing in various database applications (§ 8.1) and related notebook systems and their employed techniques for lineage tracing (§ 8.2). , and other applicable tools for saving/loading data in notebook states (§ 8.3)
8.1. PITR and Incremental Checkpointing
PITR in Relational Databases
In many DBMSs, mechanisms like ARIES (Mohan et al., 1992; Mohan and Rothermel, 1988) and its variants (Arulraj et al., 2016; Zheng et al., 2014; Soroush and Balazinska, 2013; Schule et al., 2019; Gray et al., 1981; Salem and Garcia-Molina, 1989; Verhofstad, 1978; Morrey and Grunwald, 2003) achieve durability and Point-in-time-Recovery (PITR) by combining (incremental) checkpointing and physical/operation logging (Oracle, 2024a; Group, 1996b; Soroush and Balazinska, 2013; Schule et al., 2019; Xu et al., 2017; Antonopoulos et al., 2019; Verbitski et al., 2017; Zheng et al., 2014). Persisting dirty objects (e.g., rows in relations) can enhance recovery efficiency by reducing the number of log entries to replay. ARIES (Mohan et al., 1992) identifies dirty objects at the page level by recording a RecLSN, i.e., the earliest modification time, in the dirty page table. This is possible because all the pages are controlled by the transaction system’s buffer manager. Kishu shares similarities: updates are periodically flushed. However, the core question is how to define those updates in computational notebooks. There are no buffer managers; moreover, variables/objects are dependent with inter-object references (i.e., memory pointers). Simply storing/restoring some variables independently thus can invalidate the state. Kishu addresses this with Co-Variables, atomic units of persistence that can correctly preserve inter-object dependencies. Kishu can still benefit from other PITR optimizations, e.g., non-blocking checkpointing/restoration (Mohan et al., 1992) and cost-based deterministic replay (Netezza, 2024) (§ 7.5.2), which we leave as future work.
PITR in Blob Storages
Blob storages have incorporated PITR for recovering its state to earlier points (Netezza, 2024; Group, 1996a; Azure, 2024a, b; Amazon, 2024b, a; LITEREPLICA, 2024). These systems typically log timestamped updates (e.g., Azure’s change feeds (Azure, 2024a) and SQLite’s binary log (LITEREPLICA, 2024)), which can be used to return the storage state to a previous snapshot. Computational notebooks require different approaches to determining state deltas, and Kishu addresses the unique challenge through Co-Variables and related techniques, which is the core novelty of this work.
Incremental and Differential Checkpointing in HPC
HPC and stream processing systems implement efficient incremental checkpointing (Lin et al., 2020; Keller and Bautista-Gomez, 2019; Bronevetsky et al., 2008; Apache, 2024). These works focus on high-frequency (e.g., ¿1M/second) updates to relatively simple data structures (e.g., KV-stores (Lin et al., 2020; Apache, 2024) and datasets (Keller and Bautista-Gomez, 2019; Bronevetsky et al., 2008)). For enhanced efficiency, log truncation (Lin et al., 2020) and copying (Keller and Bautista-Gomez, 2019) are employed to ensure the bounded size and robustness of incremental logs. Computational notebooks require different delta detection methods as notebooks involve more complex objects, e.g., graphs, tensors, dataframes (§ 2.2). While orthogonal, the existing optimization techniques for HPC may offer significant performance benefits if the core techniques for Kishu are expanded to handle a significantly larger number of executions (e.g., scripts) compared to interactive notebooks.
8.2. Notebook Systems and Techniques
Systems for Speeding up Data Exploration
There are a variety of works for enhancing data exploration efficiency in notebook-based systems (Wu et al., 2020; Koop and Patel, 2017; Shankar et al., 2022; Lee et al., 2021; Li et al., 2023b; Bäuerle et al., 2022; Guo and Seltzer, 2012; Brachmann et al., 2019). Reactive execution engines (Wu et al., 2020; Koop and Patel, 2017; Shankar et al., 2022) track cell reruns and rerun their dependent cells reactively to enforce consistent cell outputs. Notebook recommender systems (Lee et al., 2021; Li et al., 2023b) compute next-cell recommendations based on the current workflow. Symphony (Bäuerle et al., 2022) and B2 (Wu et al., 2020) enable point-and-click interactions with ML models and dataframes, respectively, by translation to equivalent code operations. Burrito (Guo and Seltzer, 2012) and Vizier (Brachmann et al., 2019) construct and visualize a history graph for versioning multi-language (Python, R, SQL) data science code. Diff-in-the-loop (Wang et al., 2022) enables graphical comparisons of dataframes. Kishu facilitates data exploration through efficient time-traveling, and thus, is orthogonal and complementary to these works.
Notebook State Versioning
Notebook state versioning has been explored by ForkIt (Weinman et al., 2021), which performs backtracking and branching to/from states by saving entire states with pickle (Foundation, 2023f). Our work versions notebook states in a much more efficient and robust manner compared to Forkit’s approach (which is equivalent to our DumpSession baseline in § 7) through efficient delta detection and incremental checkpointing and checkout.
Lineage Tracing in Notebooks
Lineage tracing aims to capture code dependencies, i.e., accessed and updated data of cells, and has been widely used in notebook systems for a variety of downstream tasks (e.g., reactive execution) (Manne et al., 2022; Macke et al., 2020; Ahmad et al., 2022; Li et al., 2023a; Cunha et al., 2021; Wu et al., 2020; Koop and Patel, 2017; Shankar et al., 2022; Lee et al., 2021; Li et al., 2023b; Brachmann et al., 2019; Deo et al., 2022). Tracing methods can be divided into (1) static code analysis using methods like AST decomposition (Foundation, 2023a), and (2) live code instrumentation resolving variable/data references at runtime. They are relatively (1) cheap but conservative, and (2) accurate but expensive, respectively. Lineage tracers often combine static and live analysis, to mitigate conservative assumptions of static analysis (Li et al., 2023a) with dynamically computed ’ground truths’ (Macke et al., 2020; Shankar et al., 2022; Deo et al., 2022). Unfortunately, to the best of our knowledge, all of these works, except for ElasticNotebook (Li et al., 2023a), detect modifications at the variable level, incorrectly disregard shared references between variables. Kishu’s innovations in lineage tracing is efficient live analysis through (1) modeling the state at a coarser Co-variable granularity and (2) quickly pruning update candidates. This allows Kishu to achieve low false positives (§ 7.2) with low overhead (§ 7.6).
8.3. Checkpointing Notebook Objects